i have tried many times to solve this but cant find the answer
-5g+1=7+2g

Answer by Alan3354(69443)

-5g+1=7+2g
Add 5g
1 = 7 + 7g
Subtract 7
-6 = 7g
Divide by 7
-6/7 = g